# SADA: Semantic adversarial unsupervised domain adaptation for Temporal Action Localization

### v2

@@ 4.3 Ablation studies
| Local | Global | Bkg | mAP {10,30,50}% | Avg | ||
| 30.03 | 26.62 | 17.79 | 24.81 | |||
| 30.48 | 27.07 | 18.12 | 25.22 | |||
| 30.34 | 26.66 | 17.04 | 24.68 | |||
| 29.76 | 26.63 | 17.52 | 24.64 | |||
| 31.36 | 28.01 | 18.67 | 26.01 | |||
| 30.09 | 26.88 | 17.43 | 24.80 | |||
| 32.69 | 29.17 | 19.72 | 27.19 | |||
[…]
| Local | Global | Bkg | mAP {10,30,50}% | Avg | ||
| 30.74 | 27.23 | 18.07 | 25.34 | |||
| 30.87 | 27.45 | 18.29 | 25.97 | |||
| 30.97 | 27.64 | 18.39 | 25.67 | |||
| 30.36 | 26.82 | 18.01 | 25.06 | |||
| 30.82 | 27.57 | 18.30 | 25.56 | |||
| 30.30 | 26.98 | 17.85 | 25.04 | |||
| 31.60 | 28.22 | 18.98 | 26.27 | |||
Ablation of the effect of the hyper parameters. As described in Eq. 9, we define our SADA loss using a set of hyperparameters , where each of the parameters controls the influence of a given resolution level in the overall loss. Given the relevance of these parameters, below we ablate over different variations to showcase the effect that they have in the final performance of the model.
Concretely, we define the three following scenarios:
1.
All levels to 1: This scenario sets all the parameters , keeping the contribution of each of the levels the same.
2.
First 3 levels 1: This scenario sets the first half of the levels to 1 and the rest to 0. This is and .
3.
Last 3 levels 1: This scenario sets the first half of the levels to 0 and the rest to 1. This is and .
All in all, this study attempts to clarify how sensitive is our model to this design decision. In this regard, these results (see Table V) indicate a mild sensitivity to this choice. Specifically, we observe that the second best-performing method attains an absolute difference over the optimal choice of only mAP and mAP for the two considered scenarios S1 and S3, respectively. Moreover, we observe that the worst performing naive strategy still outperforms in both scenarios 4 out of the 6 tested UDA baselines.
[…]
| Strategy | mAP {10,30,50}% | Avg | ||
| All levs 1 | 32.42 | 29.59 | 19.34 | 27.11 |
| First 3 levs 1 | 30.96 | 27.48 | 18.29 | 25.58 |
| Last 3 levs 1 | 30.54 | 27.19 | 18.28 | 25.34 |
| ”Optimal choice” | 32.69 | 29.17 | 19.72 | 27.19 |
[…]
| Strategy | mAP {10,30,50}% | Avg | ||
| All levs 1 | 30.60 | 27.36 | 18.41 | 25.45 |
| First 3 levs 1 | 31.05 | 27.64 | 18.30 | 25.66 |
| Last 3 levs 1 | 30.95 | 27.59 | 18.56 | 25.70 |
| ”Optimal choice” | 31.60 | 28.22 | 18.98 | 26.27 |
Study of the class embedding. One of the main contributions of our work is to adversarially align distributions in a class-wise fashion. Intuitively, this requires that our level-wise domain discriminator knows the class distribution that it is aligning. In this regard, in Sec. 3.5 we propose to concatenate to every anchor a learnable embeddings of its corresponding class (see Eq. 7). In this ablation, we empirically justify our choice. For this, we compare our approach against several non-learnable alternatives to encode a given class : a naive one-hot encoding, a random class-wise dense initialization, and the sinusoidal encoding originally proposed by [33].
In Table VI we show the experimental results of each of the variants tested in S1 and S3. These indicate that a naive one-hot encoding of a class is the best-performing non-learnable strategy, obtaining considerable improvements over the other two tested non-learnable baselines. This improvement is especially prominent in the first scenario, boosting the performance over the other non-learnable strategies by mAP and mAP, respectively. Moreover, to our surprise, the popular Sinusoidal encoding [33] proves to be the worst-performing method, being consistently outperformed by even the random dense encoding. Finally, we highlight that in both scenarios the use of a learnable class embedding yields the best results, improving by mAP and mAP the performance of the one-hot encoding strategy, which justifies our choice.
[…]
| Strategy | mAP {10,20,30,40,50}% | Avg | ||||
| One-hot | 32.29 | 31.09 | 28.58 | 24.88 | 19.56 | 27.28 |
| Random emb | 30.71 | 29.50 | 27.43 | 23.78 | 18.52 | 25.99 |
| Sinusoidal [33] | 29.93 | 28.57 | 26.44 | 22.87 | 17.01 | 24.96 |
| Learnable | 32.69 | 31.49 | 29.17 | 25.51 | 19.72 | 27.72 |
[…]
| Strategy | mAP {10,20,30,40,50}% | Avg | ||||
| One-hot | 31.31 | 29.97 | 27.84 | 24.28 | 18.81 | 26.44 |
| Random emb | 31.38 | 30.01 | 27.94 | 24.13 | 18.70 | 26.43 |
| Sinusoidal [33] | 31.10 | 29.78 | 27.71 | 23.99 | 18.56 | 26.23 |
| Learnable | 31.68 | 30.32 | 28.37 | 24.55 | 19.09 | 26.80 |
Ablation of the effect of the negative anchors. In this work we focus on applying domain adaptation over anchor-based localization methods, the current state-of-the-art architectures on datasets like EpicKitchens100 [66]. The success of these architectures is normally explained by the generation of a very considerable number of anchors, largely surpassing the number of actual ground-truth labels to predict. As described in Sec. 3 these are the excess anchors –i.e., negative anchors– that are not directly matched with a GT segment, consequently training the model to predict class 0 instead. In other words, in these cases we hope for the model to learn to identify these anchor embeddings as background, not predicting any action class during inference. This, however, is a cumbersome issue that most of the anchor-based methods face, which is exacerbated when doing domain adaptation in this complicated setup. In this case, it is not clear the amount of noise that these embeddings induce in the alignment strategy. Even though we leave a more thorough analysis of this issue as future work, this ablation aims to provide minimal proof of the potential negative effect that this has on the model’s performance.
[…]
| Filtered % | mAP {10,20,30,40,50}% | Avg % | ||||
| 0% | 32.69 | 31.49 | 29.17 | 25.51 | 19.72 | 27.72 |
| 25% | 32.54 | 31.87 | 30.06 | 26.99 | 21.91 | 28.67 |
| 50% | 34.46 | 33.55 | 31.29 | 27.59 | 21.18 | 29.61 |
| 75% | 35.09 | 34.53 | 32.97 | 30.05 | 23.97 | 31.32 |
| 100% | 35.57 | 35.24 | 34.16 | 30.72 | 24.50 | 32.04 |
[…]
| Filtered % | mAP {10,20,30,40,50}% | Avg % | ||||
| 0% | 31.60 | 30.29 | 28.22 | 24.47 | 18.98 | 26.71 |
| 25% | 32.76 | 31.60 | 29.81 | 26.14 | 20.24 | 28.11 |
| 50% | 35.25 | 34.25 | 32.68 | 29.25 | 23.01 | 30.89 |
| 75% | 38.39 | 37.77 | 36.40 | 33.19 | 26.33 | 34.42 |
| 100% | 43.93 | 43.59 | 42.61 | 39.26 | 31.85 | 40.25 |
[…]
| Filtered % | mAP {10,20,30,40,50}% | Avg | ||||
| 0% | 32.69 | 31.49 | 29.17 | 25.51 | 19.72 | 27.72 |
| 25% | 31.48 | 30.34 | 27.97 | 24.07 | 18.93 | 26.56 |
| 50% | 32.11 | 31.11 | 29.25 | 25.85 | 20.72 | 27.81 |
| 75% | 35.99 | 35.28 | 33.48 | 29.90 | 23.64 | 31.66 |
| 100% | 37.06 | 35.73 | 32.34 | 26.45 | 33.82 | 47.08 |
[…]
| Filtered % | mAP {10,20,30,40,50}% | Avg | ||||
| 0% | 31.60 | 30.29 | 28.22 | 24.47 | 18.98 | 26.71 |
| 25% | 32.25 | 30.46 | 28.04 | 24.54 | 19.03 | 26.86 |
| 50% | 33.43 | 32.06 | 29.98 | 26.44 | 20.55 | 28.49 |
| 75% | 38.26 | 37.33 | 35.72 | 32.53 | 25.61 | 33.89 |
| 100% | 47.08 | 46.69 | 45.65 | 42.31 | 34.42 | 43.23 |
Concretely, we propose an experiment where we progressively mask out the set of negative anchors. For this, in Table VII we randomly mask out these only during inference. Put differently, we leave all the training routines unchanged – thus considering all the anchors – but ensure that a percentage of the negative anchors are not considered during inference. This measures the predicting potential of a model that is able to perfectly deal with these embeddings. As observed in Table VII, in both considered scenarios we observe a linear performance increase, reaching an absolute improvement of mAP and mAP of the version with filtered negative anchors over the version. This is, the variant where only the positive anchors are considered over that containing both positive and all the negative ones.
In Table VIII we propose an alternative analysis. In this case, we do not only mask the negative anchors during inference but also during training. In this case, for instance, masking out of the negative anchors additionally implies not training any of the task losses on them, nor using them on our alignment loss. As observed in the results, masking them out entirely – i.e., – has an even bigger impact, reaching an absolute improvement of mAP and mAP, respectively, when comparing it to the filtering version. Nevertheless, in this case, we observe that the improvement is no longer linear, and even worsens the performance for small percentages of masking. We hypothesize this is induced by a reduction of the training data of background class. Thus, while masking some negative embeddings has a positive effect, we must also consider the negative effect that having less training data has on the ones that we do consider. Hence, the improvement in performance remains little to non-existent until this trade-off is dominated by the positive effect of masking out the majority of negative anchors. In conclusion, these experiments emphasize the need for devising carefully designed models that can properly cope with this intrinsic limitations of anchor-based methods.
Per-class analysis. Finally, we provide a more detailed class-wise analysis of the performance of SADA. Concretely, in Table IX we show the respective class-wise mAP scores of the 10 considered classes on both S1 and S3. For the analysis, we also include the results obtained with the source-only variant of our model as well as DANN [3] baseline.
[…]
| take | put | wash | open | close | insert | turn-on | cut | turn-off | pour | |
[…]
| take | put | wash | open | close | insert | turn-on | cut | turn-off | pour | |
Observe that in S1 (see Table IX bottom) our model attains the best class-wise performance in 5 of the classes, while DANN [3] does so on 3, and the source-only model on 2. In contrast, in S3 (Table IX top) our method obtains a much clearer improvement over the chosen baselines, yielding the best results in 8 of the 10 classes. Overall, we can observe that our method performs very well in the 3 majority classes of both scenarios, where we highlight the absolute improvement of SADA over DANN [3] of mAP for the class put in S3. Moreover, our method fails to improve the performance of action pour in both scenarios, which we attribute to the lack of sufficient data to operate on our proposed methodology. We also highlight that S1 presents a more challenging setup, degrading the performance in other actions such as open, close, insert or cut. This indicates the existence of intrinsic qualitative aspects that harden the adaptation when dealing with old videos. This is not the case in scenario S3, as aside from the aforementioned pour segments, it only fails to achieve the best results for the turn-on action.
4.4 Qualitative analysis
TSNE alignment study. In this study we first perform a qualitative per-class analysis of the effect of our proposed alignment loss SADA, visually depicting the level of alignment between Source and Target domain embeddings. To this end, in Fig. 5 we show the TSNE plots of the domain-invariant embeddings of the 3 majority classes (first 3 columns) and background embeddings (last column) when trained with SADA and when learning source-only. Observe that source-only (top row) yields clearly unaligned distributions with scarce to no overlap in the projected space. Our method, in contrast, presents a considerable distribution mix improving the alignment of class-wise distributions across domains. Given the anchor-based nature of our method, we have numerous background anchors –i.e., not assigned to any true action label. As observed in the last column, these are also aligned by our method (resulting from Eq. 8) effectively aligning the entire data distributions but in a semantically sensitive way.
In Fig. 5 we only include the visualizations of the embeddings that correspond to the level 0 of our multi-resolution architecture. It remains unclear, however, whether this alignment behavior remains consistent across the different resolution levels. To showcase this behavior we include in Fig. 6 - Fig. 9 the TSNE plots of the 3 majority classes and of background embeddings, respectively. Each of these figures includes the different plots at each of the first 3 resolution levels, where level 0 is the shallowest, and level 3 is the deepest studied level. We highlight that deeper levels are not meaningful to plot given the scarce number of resulting domain-invariant embeddings, caused by the downsampling that each of the levels of the architecture performs.
Observe that the influence of the alignment loss follows a similar pattern in all the resolutions. Concretely, in the three presented, we observe a considerable mixing between the source and target distributions (red and blue, respectively). This contrasts with the very clear disentanglement of the representations of both domains in the source-only version, where these present little to no overlap. We highlight what seems to be the only exception which is resolution level 3 of class 1. This one shows little improvement in the mixing of the distributions compared to the source-only variant. We also find that our overall improvement of the mixing of the distributions is consistent when analyzing the background class embeddings. Observe in Fig. 9 that in all 3 studied resolution levels, SADA improves very considerably the alignment, pushing the feature space of both domains to be domain invariant. This emphasizes the positive influence of the background alignment term of our loss (see Eq. 8).
Segment visualization. We complement our qualitative study by depicting in Fig. 10 a segment visualization of S3 (left) and one of our proposed setup for CharadesEgo (right). More specifically, in the segment visualization of S3 we can observe that the Actionformer [21] misses many of the segments in the shown video clip ignoring all the take actions and mistaking a close action for a take. Tridet [10] performs better but misses all the take actions in the first half of the video while worsening the boundary prediction of the last close. SADA improves Tridet by predicting the second take action and considerably improving the boundary of the last close action.
Similarly, note that the segment visualization on CharadesEgo (right) shows a much more complex setup, where there are many overlapping actions. For the depicted example, we must point out that the worst-performing model is the ActionFormer [21]. This model misses the action walk and one of the actions put. It also has important limitations in terms of localization, especially for the action labels drink and hold. Then there is Tridet [10], which improves ActionFormer considerably but still misses the walk action and is unable to localize the longest put action. It also predicts an action take that does not correspond to any ground truth. Instead, our method produces much more accurate results, does not miss any of the ground truth actions, and in most cases accurately localizes most of the segments. Yet it confuses the two hold actions, which is a consistent failure across all tested baselines.
5 Conclusions and limitations
In this work, we deal for the first time with Unsupervised Domain Adaptation on realistic Temporal Action Localization (TAL) scenarios. We propose a novel semantic adversarial loss that enables a more fine-grained distribution alignment compared to existing global-distribution-based approaches. Given the lack of suitable evaluation setups for this scenario, we propose a suite of 7 different benchmarks that provide a comprehensive assessment of the model performance across various domain shifts. These experiments indicate that our model yields a considerable improvement over state-of-the-art methods, which we support with extensive quantitative and qualitative results.
Limitations and future work. Despite the virtues of our proposed benchmarking scenarios, we leave as future work to complement this evaluation with long-tail scenarios, analyzing the effect of domain adaptation optimization in the presence of highly unbalanced data in terms of labels. We also leave as future work the analysis and possible benefits of the adaptation of our semantic adversarial loss to other domain adaptation problems, even to those not necessarily related to video tasks.
@@ 5 Conclusions and limitations
In this supplementary material, we first provide the implementation details needed to reproduce our work (see Sec. A). We then extend the description of Sec. IV.A of the main paper regarding our 7 proposed benchmarking scenarios (see Sec. B). In Sec. C we describe in detail our chosen UDA baselines. Finally we extend some of the provided ablation studies from Sec. IV.C (see Sec. D.1 and Sec. D.2).
Appendix A Implementation details
@@
Architecturally, all our models define the main feature extractor as a 6-level SGP feature pyramid [10] that follows a max pooling strategy and uses a downsampling rate of and an internal feature dimensionality of . We model the classification and localization heads as a sequence of three 1D-CNNs with a kernel size of . These heads are shared across embeddings of different resolution levels. Finally, we introduce level-wise domain discriminators designed as a multi-layer perceptron (MLP) of depth of width . Find in Table S1 the other relevant scenario-dependent hyperparameters related to our main contribution, the SADA loss. The rest of the hyperparameters follow the original implementation of Tridet [10].
@@
Appendix B Extended benchmark description
In this section, we describe in more depth the 7 different benchmarking scenarios that we propose in Sec. IV.A of the main text. Concretely, in Sec. B.1 we expand on the 6 different setups that build on the EpicKitchens100 dataset. For this, in Sec. B.1 we first briefly describe EpicKitchens100 [14], the base of all our benchmarks. We then illustrate in Sec. B.1 and Sec. B.1 the 2 different domain shifts that we identify, which we complement with a more detailed description of the specific scenarios derived from each of them. In Sec. B.1 we also include a detailed set of statistics for each of the proposed setups. We proceed similarly for the scenario based on Charades-Ego (see Sec. B.2). This includes a brief description of the dataset and a detailed explanation of the domain adaptation benchmark that we propose in this work based on this dataset.
@@ B.1 EpicKitchens100 scenarios
Description: EpicKitchens100 [14] is a large-scale ego-centric dataset that has recently gained a lot of attraction in the Temporal Action Localization community to test for challenging detection scenarios. This dataset extends the previous version of the dataset, namely EpicKitchens55 [66]. This dataset contains videos of 32 subjects performing daily cooking activities – e.g., wash, put, open, close – in different kitchen environments. This results in 100 hours of non-scripted videos containing 20M frames annotated with 90K action segments.
One of the key challenges of this dataset is the presence of 97 different actions, forming a very long tail action distribution. This inherent characteristic is however an issue that falls beyond the scope of this paper, given the important additional challenges that this poses in existing domain adaptation-based methods. For this reason, in all our experiments we consider only the 10 majority classes – i.e., take, put, wash, open, close, insert, turn-on, cut, turn-off and pour. As shown in Fig. S1 this pruning keeps most of the labeled action segments of the original dataset, concretely keeping up to of it (marked in blue in the histogram). Moreover, we find that this design decision does not diminish in a relevant way the challenging nature of this task. Proof of this is that the currently best-performing methods on this dataset – i.e., Actionformer [21] and Tridet [10]– still attained mAP metrics inferior to in all our proposed scenarios.
@@
Changes in the acquisition devices: The new videos that were recorded during the extension relied on a newer camera device. Importantly, this camera incorporates camera stabilization techniques. Fig. S2 visually depicts the improvement in the stability of new over old videos.
@@
For this reason, we identify a clear, understandable domain gap. This is the color of the kitchen counters. This is an essential part of the background information, and thus, we hypothesize that adapting to this factor is also critical to obtain high-performing models. In this regard, we split the data into three different domains: dark, white, and other types of kitchen counters. This provides a clear domain gap that we can visualize and thus, understand (see Fig. S3). We refer interested readers to the file domain shifts from the supplementary which shows video content that further depicts the differences between domains in all the proposed scenarios.
@@
All in all, we stress that all our proposed scenarios provide sufficiently even splits to design single-domain setups where we can ensure that there is no easy generalization of the target domain by means of the source training data (as argued in Sec. IV.A of the main text).
Study of the benchmark statistics: In Sec. IV.A of the main text we propose a set of 6 new scenarios to perform UDA on egocentric videos based on EpicKitchens100 [14]. To do so we rely on two main factors to split the original data from EpicKitchens100 [14] into domains: color of the kitchen counter and year of acquisition –i.e., if the data was already part of EpicKitchens55 [66]. In this regard, in Table S2 we detail the relevant statistics according to these splits where we present the number of videos, the number of segments, the average length of these segments, and finally the class-wise number of segments to depict the overall distribution of actions.
Given this data splits, we define 6 different scenarios according to the description of Sec. IV.B. In this respect, in Table S3 and Table S4 we present the aforementioned statistics for the source and target domain of each scenario, respectively. Observe that thanks to our careful experimental setup, we are able to define scenarios with considerably large source and target domains. This contrasts with other experimental setups like that proposed by [61] which define the source and target domain of a scenario as the videos of a single kitchen, respectively. We find that this, in consequence, yields domains with only 15-28 videos, which we deem insufficient to train state-of-the-art train large architectures.

### v3

@@ 4.3 Ablation studies
Qualitative analysis. We complement our quantitative results with a qualitative study. For this, we depict in Fig. 4 a segment visualization of S3 of our proposed method versus the chosen baseline models. In this visualization, we can observe that the Actionformer misses many of the segments in the shown video clip ignoring all the take actions and mistaking a close for a take. Tridet performs better but misses all the take actions in the first half of the video while worsening the boundary prediction of the last close. SADA improves Tridet by predicting the second take and considerably improving the boundary of the last close action.
Fig. 5 also shows the TSNE plots of the domain-invariant embeddings of SADA with respect to learning source-only. Given that SADA is an adversarial-based class-wise loss, we depict the plots of the 3 majority action classes (first 3 columns). Observe that SO (top row) yields clearly unaligned distributions with scarce to no overlap in the projected space. Our method, in contrast, presents a considerable distribution mix improving the alignment of class-wise distributions across domains. Given the anchor-based nature of our method, we have numerous background anchors –i.e., not assigned to any GT label. As observed in the last column, these are also aligned by our method (see Eq. 8) therefore effectively aligning the entire data distributions but in a semantically sensitive way. See the Supp. for the complete study at different resolution levels.
5 Conclusions
In this work, we deal for the first time with Unsupervised Domain Adaptation on realistic Temporal Action Localization (TAL) scenarios. We propose a novel semantic adversarial loss that enables a more fine-grained distribution alignment compared to existing global-distribution-based approaches. Given the lack of suitable evaluation setups for this scenario, we propose a suite of 7 different benchmarks that provide a comprehensive assessment of the model performance across various domain shifts. These experiments indicate that our approach yields a considerable improvement over state-of-the-art methods, which we support with extensive quantitative and qualitative results.
Ackwoledgements: This work has been partially supported by the Spanish project PID2022-136436NB-I00 and by ICREA under the ICREA Academia programme.
@@ 5 Conclusions
In this supplementary material, we first provide the implementation details needed to reproduce our work (see Sec. A). We then extend the description of Sec. 4.1 regarding our 7 proposed benchmarking scenarios. In Sec. C we describe in detail our chosen UDA baselines. We then complement the provided ablation studies from Sec. 4.3 (see Sec. D.1 and Sec. D.2). In Sec. E we ablate over the importance of properly tunning the hyperparameter and in Sec. F we provide empirical evidences that justify the choice of using learnable class embeddings. In Sec. G we provide a comprehensive study of the per-class performance of our model on multiple scenarios. In Sec. H we ablate over the capacity of SADA to mitigate the potential negative impact of background anchors, and provide empirical evidences of the need of further research to continue alleviating the degradation in performance that they induce. Finally, in Sec. I we provide complementary qualitative results, including further segment visualizations (see Sec. I.1) and TSNE plots (see Sec. I.2).
A Implementation details
@@
Architecturally, all our models define the main feature extractor as a 6-level SGP feature pyramid [48] that follows a max pooling strategy and uses a downsampling rate of and an internal feature dimensionality of . We model the classification and localization heads as a sequence of three 1D-CNNs with a kernel size of . These heads are shared across embeddings of different resolution levels. Finally, we introduce level-wise domain discriminators designed as a multi-layer perceptron (MLP) of depth and width . Find in Tab. A the other relevant scenario-dependent hyperparameters related to our main contribution, the SADA loss. The rest of the hyperparameters follow the original implementation of Tridet [48].
@@
B Extended benchmark description
In this section we describe in more depth the 7 different benchmarking scenarios that we propose in Sec. 4.1. Concretely, in Sec. B.1 we expand on the 6 different setups that build on the EpicKitchens100 dataset. For this, in Sec. B.1 we first briefly describe EpicKitchens100 [14], the base of all our benchmarks. We then illustrate in Sec. B.1 and Sec. B.1 the 2 different domain shifts that we identify, which we complement with a more detailed description of the specific scenarios derived from each of them. In Sec. B.1 we also include a detailed set of statistics for each of the proposed setups. We proceed similarly for the scenario based on Charades-Ego (see Sec. B.2). This includes a brief description of the dataset and a detailed explanation of the domain adaptation benchmark that we propose in this work based on this dataset.
@@ B.1 EpicKitchens100 scenarios
Description: EpicKitchens100 [14] is a large-scale ego-centric dataset that has recently gained a lot of attraction in the Temporal Action Localization community to test for challenging detection scenarios. This dataset extends the previous version of the dataset, namely EpicKitchens55 [13]. This dataset contains videos of 32 subjects performing daily cooking activities – e.g., wash, put, open or close – in different kitchen environments. This results in 100 hours of non-scripted videos containing 20M frames annotated with 90K action segments.
One of the key challenges of this dataset is the presence of 97 different actions, forming a very long tail action distribution. This inherent characteristic is however an issue that falls beyond the scope of this paper, given the important additional challenges that this poses in existing domain adaptation-based methods. For this reason, in all our experiments we consider only the 10 majority classes – i.e., take, put, wash, open, close, insert, turn-on, cut, turn-off and pour. As shown in Fig. A this pruning keeps most of the labeled action segments of the original dataset, concretely keeping up to of it (marked in blue in the histogram). Moreover, we find that this design decision does not diminish in a relevant way the challenging nature of this task. Proof of this is that the currently best-performing methods on this dataset – i.e., Actionformer [63] and Tridet [48]– still attained mAP metrics inferior to in all our proposed scenarios.
@@
Changes in the acquisition devices: The new videos that were recorded during the extension relied on a newer camera device. Importantly, this camera incorporates camera stabilization techniques. Fig. B visually depicts the improvement in the stability of new over old videos.
@@
For this reason, we identify a clear, understandable domain gap. This is the color of the kitchen counters. This is an essential part of the background information, and thus, we hypothesize that adapting to this factor is also critical to obtain high-performing models. In this regard, we split the data into three different domains: dark, white, and other types of kitchen counters. This provides a clear domain gap that we can visualize and thus, understand (see Fig. C).
@@
All in all, we stress that all our proposed scenarios provide sufficiently even splits to design single-domain setups where we can ensure that there is no easy generalization of the target domain by means of the source training data (as argued in Sec. 4.1).
Study of the benchmark statistics: In Sec. 4.1 we propose a set of 6 new scenarios to perform UDA on egocentric videos based on EpicKitchens100 [14]. To do so we rely on two main factors to split the original data from EpicKitchens100 [14] into domains: color of the kitchen counter and year of acquisition –i.e., if the data was already part of EpicKitchens55 [13]. In this regard, in Tab. B we detail the relevant statistics according to these splits where we present the number of videos, the number of segments, the average length of these segments, and finally the class-wise number of segments to depict the overall distribution of actions.
Given this data splits, we define 6 different scenarios according to the description of Sec. 4.2. In this respect, in Tab. C and Tab. D we present the aforementioned statistics for the source and target domain of each scenario, respectively. Observe that thanks to our careful experimental setup, we are able to define scenarios with considerably large source and target domains. This contrasts with other experimental setups like that proposed by [57] which define the source and target domain of a scenario as the videos of a single kitchen, respectively. We find that this, in consequence, yields domains with only 15-28 videos, which we deem insufficient to train state-of-the-art large architectures.